OpenStudy (anonymous):

can some one please explain.. |8x + 8| + 3 = 35

OpenStudy (jdoe0001):

in |8x + 8| + 3 = 35 absolute value equations, you have to bear in mind that, what they're really saying is +1(8x + 8) + 3 = 35 one answer for "x" -1(8x + 8) + 3 = 35 another answer for "x" too so, to get the VALUES of "x", you solve both of those equations

OpenStudy (anonymous):

Absolute value just means that anything value that comes out will always be positive. To solve for x, let's first move 3 on to the right side:\[\bf |8x+8|+3=35 \implies |8x+8|=32\]Now by the properties of Absolute Value, we can factor out the 8:\[\bf |8x+8|=32 \implies |8(x+1)|=32 \implies |8|*|x+1|=32 \implies |x+1|=4\]Now since the absolute value makes everything positive, we get that x + 1 = 4 or x + 1 = -4 because in either case, the answer from the absolute value would come out to be positive 4:\[\bf |x+1|=4 \implies x+1=4 \ \ or \ \ x+1=-4\]Solve for each case separately to get:\[\bf x+1= 4 \implies x = 3\]\[\bf x+1=-4 \implies x = -5\]And you're done.

OpenStudy (anonymous):

@Moonraven